OpenStudy (anonymous):

True or false and explain. 1. Lnsqrt(X+5) = 1/2 Ln X+5 2. Ln^2(X-4) = 2Ln(X-4) 3. e^-(Ln3 = -3 4.e^(Ln2)x = 2X

OpenStudy (anonymous):

1. Lnsqrt(X+5) = 1/2 Ln X+5\[\ln \sqrt{x+5}=\ln(x+5)^{1/2}\]

OpenStudy (anonymous):

\[=\frac{1}{2}\ln(x+5)\] True

OpenStudy (anonymous):

2. Ln^2(X-4)?

OpenStudy (anonymous):

2. I don't get question

OpenStudy (anonymous):

\[\ln ^{2(x-4)}\]?

OpenStudy (anonymous):

close! but the X-4 is not in the exponent. just bring it down.

OpenStudy (anonymous):

\[Ln^2(X-4)\]